Write an equation of the line containing the given point and parallel to the given line
(3,-3) 7x-2y=5
Please show me how to get the correct answer thank you

Answer by richwmiller(17219)

the slope formula is
y=mx+b
We want this equation in that form to calculate slope
7x-2y=5
Subtract 7x from both sides
-2y=5-7x
divide both sides by -2
y=5/-2-7x/-2
minus over minus = plus
y=-2.5+3.5x
rearrange
y=3.5x-2.5
m=3.5 and y intercept=-2.5
parallel lines have the same slope
so now back to the task
y=mx+b
y=3.5x+b
insert the point coordinates and solve for b
-3=3.5*(3)+b
-3=10.5+b
b=-13.5
y=3.5x-13.5